# Thermodynamic Cost of Regeneration in a Quantum Stirling Cycle

### v1

@@ Abstract
We study the regenerative quantum Stirling heat engine cycle within the standard weak-coupling, Markovian open quantum system framework. We point out that the regeneration process is not thermodynamically free in a reduced open-system description, and we treat the required work input as an explicit regeneration cost by modifying the cycle efficiency accordingly. We consider two working substances—a single spin-$1/2$ and a pair of interacting spin-$1/2$ particles—and investigate the cycle performance by taking the regeneration cost at its minimum value set by the Carnot heat-pump limit. For comparison, we also analyze the conventional Stirling cycle without regeneration under the same conditions. The super-Carnot efficiencies reported under the cost-free regeneration assumption disappear once the regeneration cost is included: the modified efficiency stays below the Carnot bound, while still remaining higher than the efficiency of the conventional Stirling cycle. For the conventional Stirling cycle, we provide a rigorous Carnot bound using quantum relative entropy, whereas for the regenerative cycle we derive a sufficient lower bound on the regeneration cost that guarantees thermodynamic consistency. Finally, we suggest three candidate quantum regenerator models for future work.
@@ 1 Introduction
In this study, departing from the standard cost-free passive heat buffer assumption, we focus on the functioning of the regenerator and account for a finite thermodynamic cost associated with heat regeneration. Our motivation is as follows. In the four-stroke formulation of a regenerative quantum cycle, it is typically assumed that, at the end of each regenerative isochoric stroke, the working substance—after interacting with the regenerator—thermalizes to the temperature of the subsequent isothermal bath (i.e., the cold bath at $T_{c}$ after the first isochore and the hot bath at $T_{h}$ after the second), so that its reduced state can be represented by a Gibbs (thermal) state [2, 7]. Within the standard weak coupling, Markovian open quantum system framework, this approach effectively models the regenerator as a high heat capacity reservoir with an effective temperature, i.e., as an effective heat bath [57, 58]. Consequently, the regenerator plays the role of a cold bath during the first regenerative isochoric stroke and that of a hot bath during the second. For this reason, regeneration requires that heat stored at a low temperature level be made usable at a higher temperature level. In other words, upgrading energy/heat available at the $T_{c}$ level to the $T_{h}$ level is, from a second-law perspective, a heat-pumping (heat-upgrading) problem: it cannot occur spontaneously and therefore must be driven by an external source of work.
@@ 2 Regenerative Quantum Stirling Cycle
As with the most extensively analyzed Otto and Carnot cycles, the Stirling cycle has also been generalized to the quantum domain in many studies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In direct analogy with its classical counterpart, the quantum Stirling cycle consists of four strokes - two quantum isochoric and two quantum isothermal processes - as illustrated in Fig. 1. The key point in the quantum construction is that, during all strokes of the cycle, the working substance is in thermal contact with external thermal reservoirs [5]. This situation is described within the standard open quantum systems framework, where the system-bath coupling is assumed to be weak and the Born-Markov and secular (rotating-wave) approximations are satisfied [57, 58]. Under these conditions, the dynamics of the working substance can be characterized by its reduced density matrix $\rho(t)$, which evolves according to a Lindblad-form quantum master equation, $\dot{\rho}(t)=-i/\hbar\big[H(\lambda(t)),\rho(t)\big]+\mathcal{L}_{t}[\rho(t)]$. Here, $\hbar$ is the reduced Planck constant, $H(\lambda(t))$ is the Hamiltonian of the working substance, and $\lambda(t)$ denotes a time-dependent control (tuning) parameter that is externally driven. The dissipator $\mathcal{L}_{t}[\rho(t)]$ is a superoperator that encodes the irreversible thermal action of the reservoir on the system [57, 58]. Within this framework, the two quantum isothermal strokes are modeled as quasistatic processes, i.e., $\mathrm{d}\lambda(t)/\mathrm{d}t\rightarrow 0$, so that the state remains (approximately) Gibbsian at the bath temperature at each instant [5]. This requires that the driving is much slower than the bath-induced relaxation. In contrast, during the isochoric strokes the control parameter $\lambda$ is kept fixed, and the role of the stroke is purely thermalization: by the end of the isochoric process (i.e., in the steady state), the working substance is assumed to have relaxed to a Gibbs (canonical) state at the temperature of the relevant heat bath [5]. The cycle processes are detailed below.
@@
In previous studies [7, 8, 9, 10, 11, 12, 2], the efficiency has been defined in a standard way as the ratio of the net work output to the heat absorbed from the hot reservoir, i.e., $\eta=W/Q_{h}$. Refs. [7, 10, 23] have shown - and we will likewise demonstrate in the following analysis - that the efficiency can surpass the classical Carnot limit defined solely by the temperatures of the external reservoirs, $\eta_{C}=1-T_{c}/T_{h}$. It was noted that this does not violate the second law, since in these regimes the regenerator draws additional energy in the form of heat from the hot bath [7, 2].
@@
In this study, our main claim is as follows. Based on the above construction of the cycle and on the open quantum system treatment of weak system-bath interactions, there exists an unavoidable hidden thermodynamic cost that cannot be neglected and should be included in the definition of efficiency. In the thermodynamic cycle analysis formulated in terms of the reduced dynamics of the working substance, the regenerator is effectively treated as a heat bath at temperature $T_{c}$ or $T_{h}$ during stroke (2) or (4), respectively. In this formulation, it is implicitly implied that the regeneration process requires the heat $|Q_{2}|$ stored at the lower temperature level $T_{c}$ to be usable in a practical way at the higher temperature level $T_{h}$. This heat-upgrading process corresponds to pumping heat from a colder side to a hotter one, and by the second law of thermodynamics, it cannot occur without an external work input [56]; therefore, regeneration effectively entails an additional work requirement to make $|Q_{2}|$ usable at the higher temperature level $T_{h}$. In this work, we treat the required work input as a regeneration cost, and we denote it by $W_{\mathrm{cost}}$. The most conservative estimate for $W_{\mathrm{cost}}$ is obtained from the ideal Carnot heat-pump (refrigerator) limit. Because this corresponds to a reversible process (no entropy production), it gives the minimum possible work: $W_{\mathrm{cost}}=|Q_{2}|\,((T_{h}-T_{c})/T_{c})$, where $|Q_{2}|$ denotes the magnitude of the heat initially delivered to the regenerator by the working substance.
@@
In the present case, if the cost is treated purely as an energetic input and the cycle efficiency is defined as a benefit-to-input ratio, then $W_{\mathrm{cost}}$ must be counted as part of the input energy [29, 43]. Accordingly, we define a modified efficiency for the regenerative quantum Stirling cycle as the work produced per unit total resource consumption, which can be written as
@@
Here $Q_{h}$ is given in Eq. (2) and $W_{\mathrm{cost}}=|Q_{2}|\,((T_{h}-T_{c})/T_{c})$. This modified efficiency definition constitutes a central result of the present study. In particular, it will be applied to two previously studied examples of regenerative Stirling cycles which were reported to operate beyond the Carnot efficiency [7], in order to reassess their performance within a thermodynamically consistent framework.
@@
For the regenerative Stirling cycle, the situation is more subtle. If the regenerative work cost $W_{\mathrm{cost}}$ is neglected, the efficiency can exceed the Carnot limit, as already reported in Refs. [7, 10, 23] and as shown in Figs. 2 and 3. When $W_{\mathrm{cost}}$ is included in the efficiency definition (Eq. (3)) and chosen at its lower bound given by the Carnot heat pump limit, the second law is numerically upheld in all the considered parameter regimes. However, in contrast to Eq. (4) as derived for the conventional cycle, this phenomenological description does not allow us to prove analytically that the Carnot efficiency is a strict upper bound; that is, the deviation $\eta_{C}-\eta$ cannot be written in a closed form that is manifestly positive.
There are two main reasons why the second law cannot be proven analytically for the proposed cycle model, even after including $W_{\mathrm{cost}}$. First, $W_{\mathrm{cost}}$ does not have a unique, precise expression, and there is no rigorous theory that tells us how to incorporate it consistently into a thermodynamic cycle defined only on the reduced state of the working substance [4, 40, 41, 42, 43]. In our work, we choose $W_{\mathrm{cost}}$ to be its lower bound determined by the Carnot heat-pump limit, but in realistic implementations one typically has $W_{\mathrm{cost}}>|Q_{2}|(T_{h}-T_{c})/T_{c}$, primarily due to internal irreversibilities and other possible cost contributions, with the actual value depending on the physical design of the regenerator. Second, the regenerator is not modeled as an active quantum component of the cycle, but is instead treated as an effective classical reservoir, and its influence is included only through an additional penalty term that accounts for internal heat-transfer cost. Taken individually or together, these two features — the design-dependent nature of $W_{\mathrm{cost}}$ and the purely phenomenological treatment of the regenerator — imply that the irreversibilities of the regenerative isochoric processes cannot be collected into a compact, relative-entropy-like expression. Consequently, one cannot derive a clear, always-positive ”Carnot-deficit” for the regenerative cycle analogous to Eq. (4). Instead, for completeness, we can derive a sufficient lower bound for the regeneration cost that ensures the “Carnot deficit” is always non-negative, i.e., $\eta_{C}-\eta\geq 0$. Using Eq. (4) to derive the work output together with the modified efficiency definition (3), one can formally rewrite $\eta_{C}-\eta$ and extract a sufficient lower bound for the regeneration cost (see Appendix B for the calculation details):
This bound guarantees $\eta_{C}-\eta\geq 0$, but it depends on the system parameters, in particular on the entropy production term $\Sigma=S(\rho_{B}\|\rho_{C})+S(\rho_{D}\|\rho_{A})$ defined for the conventional Stirling cycle, and on the heat exchanges $Q_{i}$, and therefore has no simple relation to the Carnot heat-pump cost $|Q_{2}|(T_{h}-T_{c})/T_{c}$. In realistic implementations, the regeneration cost must satisfy both the thermodynamic constraint $W_{\mathrm{cost}}\geq|Q_{2}|(T_{h}-T_{c})/T_{c}$ and the sufficient condition in Eq. (5). In other words, the second law imposes a more stricter condition on the cost term, namely,
@@
As noted above, for a physically realistic proposal of a regenerative quantum Stirling cycle, the regenerator itself should be modeled as a quantum system that actively participates in the isochoric strokes; in that case, the system–regenerator coupling and correlations become essential. Three candidate quantum-regenerator models suggest themselves. First, the regenerator could be modeled not as a Markovian heat bath but as a non-Markovian reservoir [41, 62, 63]. Unlike memoryless Markovian baths, non-Markovian reservoirs do not necessarily cause irreversible information loss and can exhibit information and energy back-flow [63], allowing a portion of the heat absorbed at stage (2) to be returned. Second, the regenerator can be modeled as an auxiliary quantum system on the same footing as the working substance [5, 47, 64, 65, 66, 54]. The heat rejected during the stroke (2) can induce eigenstate transitions in this auxiliary system, thereby storing energy in its populations. Algorithmic or adiabatic control protocols can subsequently be applied to reconvert the stored energy and regenerate the heat during the stage (4). Third, the regenerator may be described within a repeated–interaction (collision–model) framework, in which the working medium successively interacts with a finite chain of ancillary quantum systems that play the role of a structured reservoir with memory [67, 68, 38, 39]. During stage (2), a subset of these ancillas absorbs energy from the working medium, while during stage (4) the same ancillas are brought back into contact with the working medium and partially release the stored energy, thereby allowing partial heat regeneration. In all of the above proposals, the quantum modeling of the regenerator is intrinsically tied to the specific structure of the working substance; therefore, it is generally not possible to formulate a universal, model-independent quantum-regenerator description. Moreover, the isochoric strokes can terminate in non-equilibrium states. Consequently, additional thermalization steps are required prior to the subsequent isothermal contacts, so that the full cycle can be naturally described as a modified six-stroke regenerative Stirling cycle. A detailed investigation and their mathematical formulations of these possibilities are left for future work.

### v2

@@ Abstract
We study the standard four-stroke regenerative quantum Stirling heat engine cycle, which assumes local thermal equilibrium at each stage, within the standard weak-coupling, Markovian open quantum system framework. We point out that the regeneration process is not thermodynamically free in a reduced open-system description, and we treat the required work input as an explicit regeneration cost by modifying the cycle efficiency accordingly. We consider two working substances—a single spin-$1/2$ and a pair of interacting spin-$1/2$ particles—and investigate the cycle performance by taking the regeneration cost at its minimum value set by the Carnot heat-pump limit. For comparison, we also analyze the conventional Stirling cycle without regeneration under the same conditions. The super-Carnot efficiencies reported under the cost-free regeneration assumption disappear once the regeneration cost is included: the modified efficiency stays below the Carnot bound, while still remaining higher than the efficiency of the conventional Stirling cycle. For the conventional Stirling cycle, we provide a rigorous Carnot bound using quantum relative entropy, whereas for the regenerative cycle we derive a sufficient lower bound on the regeneration cost that guarantees thermodynamic consistency. Finally, we suggest three candidate quantum regenerator models for future work.
@@ 1 Introduction
In this study, we revisit the standard four-stroke regenerative quantum Stirling cycle, which assumes local thermal equilibrium at each stage [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Departing from the standard cost-free passive heat buffer assumption, we focus on the functioning of the regenerator and account for a finite thermodynamic cost associated with heat regeneration. By explicitly introducing this mandatory work input, we demonstrate that the previously reported super-Carnot efficiencies are merely artifacts of an incomplete thermodynamic accounting [7, 10, 23]. Once this hidden cost is integrated into the cycle, thermodynamic consistency is restored, and the efficiency remains strictly bounded by the Carnot limit. Our motivation is as follows: In the four-stroke formulation of a regenerative quantum cycle, it is typically assumed that, at the end of each regenerative isochoric stroke, the working substance—after interacting with the regenerator—thermalizes to the temperature of the subsequent isothermal bath (i.e., the cold bath at $T_{c}$ after the first isochore and the hot bath at $T_{h}$ after the second), so that its reduced state can be represented by a Gibbs (thermal) state [2, 7]. Within the standard weak coupling, Markovian open quantum system framework, this approach effectively models the regenerator as a high heat capacity reservoir with an effective temperature, i.e., as an effective heat bath [57, 58]. Consequently, the regenerator plays the role of a cold bath during the first regenerative isochoric stroke and that of a hot bath during the second. For this reason, regeneration requires that heat stored at a low temperature level be made usable at a higher temperature level. In other words, upgrading energy/heat available at the $T_{c}$ level to the $T_{h}$ level is, from a second-law perspective, a heat-pumping (heat-upgrading) problem: it cannot occur spontaneously and therefore must be driven by an external source of work.
@@ 2 Regenerative Quantum Stirling Cycle
As with the most extensively analyzed Otto and Carnot cycles, the Stirling cycle has also been generalized to the quantum domain in many studies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In direct analogy with its classical counterpart, the quantum Stirling cycle consists of four strokes - two quantum isochoric and two quantum isothermal processes - as illustrated in Fig. 1. The key point in the quantum construction is that, during all strokes of the cycle, the working substance is in thermal contact with external thermal reservoirs [5]. This situation is described within the standard open quantum systems framework, where the system-bath coupling is assumed to be weak and the Born-Markov and secular (rotating-wave) approximations are satisfied [57, 58]. Under these conditions, the dynamics of the working substance can be characterized by its reduced density matrix $\rho(t)$, which evolves according to a Lindblad-form quantum master equation, $\dot{\rho}(t)=-i/\hbar\big[H(\lambda(t)),\rho(t)\big]+\mathcal{L}_{t}[\rho(t)]$. Here, $\hbar$ is the reduced Planck constant, $H(\lambda(t))$ is the Hamiltonian of the working substance, and $\lambda(t)$ denotes a time-dependent control (tuning) parameter that is externally driven. The dissipator $\mathcal{L}_{t}[\rho(t)]$ is a superoperator that encodes the irreversible thermal action of the reservoir on the system [57, 58]. Within this framework, the two quantum isothermal strokes are modeled as quasistatic processes, i.e., $\mathrm{d}\lambda(t)/\mathrm{d}t\rightarrow 0$, so that the state remains (approximately) Gibbsian at the bath temperature at each instant [5]. This requires that the driving is much slower than the bath-induced relaxation. In contrast, during the isochoric strokes the control parameter $\lambda$ is kept fixed, and the role of the stroke is purely thermalization: by the end of the isochoric process (i.e., in the steady state), the working substance is assumed to have relaxed to a Gibbs (canonical) state at the temperature of the relevant heat bath [5]. The processes of the standard four-stroke regenerative cycle, which assumes local thermal equilibrium at each point, are detailed below.
@@
The framework of the quantum regenerative Stirling cycle discussed so far—defined by a four-stroke process where the working substance attains a Gibbs thermal state at the corresponding bath temperature at the end of each isochoric stroke—has recently been widely explored across various quantum working substances in the literature [7, 8, 9, 10, 11, 12, 2, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In such studies, the efficiency has been defined in a standard way as the ratio of the net work output to the heat absorbed from the hot reservoir, i.e., $\eta=W/Q_{h}$. Refs. [7, 10, 23] have shown - and we will likewise demonstrate in the following analysis - that the efficiency can surpass the classical Carnot limit defined solely by the temperatures of the external reservoirs, $\eta_{C}=1-T_{c}/T_{h}$. It is worth noting here that this apparent performance gain beyond Carnot limits actually stems from the incomplete accounting of a mandatory work input; it relies on the regenerator, which, as we will show, implicitly requires an additional energy expenditure for the regeneration process—a cost not accounted for in the standard definition of $\eta=W/Q_{h}$ [7, 2].
@@
However, a fundamental challenge arises: determining the energy expended for regeneration without any knowledge of the regenerator’s specific material structure or physical properties. Fortunately, by focusing on the thermodynamic task the regenerator performs, we can assign the minimum energy cost allowed by nature—the Carnot limit. Accordingly, our main claim is as follows. Based on the above construction of the cycle and on the open quantum system treatment of weak system-bath interactions, there exists an unavoidable hidden thermodynamic cost that cannot be neglected and should be included in the definition of efficiency. In the thermodynamic cycle analysis formulated in terms of the reduced dynamics of the working substance, the regenerator is effectively treated as a heat bath at temperature $T_{c}$ or $T_{h}$ during stroke (2) or (4), respectively. In this formulation, it is implicitly implied that the regeneration process requires the heat $|Q_{2}|$ stored at the lower temperature level $T_{c}$ to be usable in a practical way at the higher temperature level $T_{h}$. This heat-upgrading process corresponds to pumping heat from a colder side to a hotter one, and by the second law of thermodynamics, it cannot occur without an external work input [56]; therefore, regeneration effectively entails an additional work requirement to make $|Q_{2}|$ usable at the higher temperature level $T_{h}$. In this work, we treat the required work input as a regeneration cost, and we denote it by $W_{\mathrm{cost}}$.
The most conservative estimate for $W_{\mathrm{cost}}$ is obtained from the ideal Carnot refrigerator or heat-pump limit. Whether one considers extracting the heat $|Q_{2}|$ from the cold side at $T_{c}$ (refrigeration) or delivering the total conserved energy $|Q_{2}|+W_{\mathrm{cost}}$ to the hot side at $T_{h}$ (heat pumping), both perspectives necessitate the same minimum work input to make the heat $|Q_{2}|$ usable at the higher temperature level $T_{h}$, given by $W_{\mathrm{cost}}=|Q_{2}|((T_{h}-T_{c})/T_{c})$. Alternatively, this cost can be understood through the lens of exergy: the minimum energy required to upgrade an entropy unit, $\Delta S=|Q_{2}|/T_{c}$, stored at $T_{c}$ to the $T_{h}$ level is exactly $W_{\mathrm{cost}}=T_{h}\Delta S-T_{c}\Delta S=|Q_{2}|((T_{h}-T_{c})/T_{c})$. Furthermore, this $W_{\mathrm{cost}}$ perfectly restores the entropy balance during the regeneration process. If we were to transfer $|Q_{2}|$ directly from $T_{c}$ to $T_{h}$ without work, the entropy change of the effective reservoirs would be $\Delta S_{\mathrm{reg}}=-|Q_{2}|/T_{c}+|Q_{2}|/T_{h}<0$, resulting in a direct violation of the second law. Including the reversible work input ensures the zero entropy production condition, $\Delta S_{\mathrm{reg}}=-|Q_{2}|/T_{c}+(|Q_{2}|+W_{\mathrm{cost}})/T_{h}=0$. Fundamentally, these four thermodynamic approaches reveal that $W_{\mathrm{cost}}$ is an unavoidable energetic penalty, strictly required to make the heat $|Q_{2}|$ accessible to the working substance at the $T_{h}$ level during the $D\to A$ regeneration process. At the same time, all these fundamental thermodynamic approaches converge to a unique minimum cost for heat regeneration, given by:
$W_{\mathrm{cost}}=|Q_{2}|\left(\frac{T_{h}-T_{c}}{T_{c}}\right).$ (3)
It should be noted that, without assigning a specific internal structure to the regenerator and basing the analysis solely on its thermodynamic function, the Carnot heat-pump limit given by Eq. (3) is the only unique, model-independent expression that can be formulated for this energy cost. Furthermore, more realistic regenerator models will inevitably involve internal irreversibilities and additional operational costs. Consequently, for such physical implementations, the actual regeneration cost will strictly exceed this ideal bound, requiring $W_{\mathrm{cost}}>|Q_{2}|((T_{h}-T_{c})/T_{c})$.
@@
In standard thermodynamics, cycle efficiency is defined as the ratio of useful energy output to total energy input. Here, the regeneration cost, $W_{\mathrm{cost}}$, acts as external work supplied to make regeneration feasible at the $T_{h}$ level. When this cost is treated purely as an energetic input, it must be explicitly included in the total input energy [29, 43]. Accordingly, we define a modified efficiency for the regenerative quantum Stirling cycle as the net work produced per unit of total resource consumption, which can be expressed as:
@@
Here, $Q_{h}$ is given in Eq. (2) and $W_{\mathrm{cost}}=|Q_{2}|\,((T_{h}-T_{c})/T_{c})$. Equation (4) measures the global efficiency of the composite system, where the denominator represents the total energy bill paid. This total energetic cost encompasses the standard heat input from the hot reservoir, $Q_{h}$, alongside the additional energetic penalty, $W_{\mathrm{cost}}$, paid to operate the regenerative heat pump. It is also important to highlight that an alternative modified efficiency, defined as $\eta=(W-W_{\mathrm{cost}})/Q_{h}$, is frequently used in the literature, particularly in shortcut-to-adiabaticity calculations [42] to include the energetic cost of the shortcut—a protocol specifically designed to enhance the work and power output in finite-time operations. However, this definition is not appropriate for the present regenerative analysis. In our model, regeneration does not alter the net work; the system yields the exact same work output, $W$, in both regenerative and non-regenerative cycles. Furthermore, since the regenerator in our framework acts effectively as a thermal bath, the cycle duration—and consequently the power output—remains unchanged. Since regeneration is fundamentally a process aimed strictly at improving efficiency by recycling internal heat (thereby reducing the external heat demand) rather than boosting work or power, treating the energetic cost as an additional input in the denominator is the physically consistent approach. This modified efficiency definition constitutes a central result of the present study. In particular, it will be applied to two previously studied examples of regenerative Stirling cycles, which were reported to operate beyond the Carnot efficiency [7], in order to reassess their performance within a thermodynamically consistent framework.
@@
For the regenerative Stirling cycle, incorporating the regeneration cost profoundly clarifies the thermodynamic landscape. If the regenerative work cost, $W_{\mathrm{cost}}$, is neglected, the efficiency artificially exceeds the Carnot limit, as previously reported in Refs. [7, 10, 23] and illustrated in Figs. 2 and 3. However, when $W_{\mathrm{cost}}$ is rigorously included in the efficiency definition (Eq. (4)) and evaluated at its fundamental lower bound set by the Carnot heat-pump limit, the second law is strictly upheld. Yet, in contrast to the conventional Stirling cycle where the Carnot deficit ($\eta_{C}-\eta$) can be expressed in a manifestly positive closed form across all parameter regimes (as derived in Eq. (5)), the derivation for the regenerative cycle is more intricate and does not yield such a universally positive analytical expression (see Appendix B). Instead, the analytical proof depends explicitly on the operating temperature regimes. Crucially, our formalism allows us to analytically prove that the Carnot efficiency remains a definitive upper bound (i.e., $\eta_{C}-\eta>0$) for all temperature regimes where $T_{h}\geq 2T_{c}$. In the regime where $T_{h}<2T_{c}$, the Carnot deficit cannot be reduced to a strictly positive closed form. Nevertheless, the absence of a universally positive analytical proof in this specific parameter space does not signal a violation of the second law; rather, it reflects the inherent mathematical limitation of projecting a reversible macroscopic bound onto an irreversible, reduced quantum process. To decisively confirm the validity of the Carnot bound in this mathematically elusive $T_{h}<2T_{c}$ regime, we point to our exact numerical results in Figs. 2 and 3. For these evaluations, the bath temperatures are explicitly set to $T_{h}=3$ and $T_{c}=2$, perfectly capturing the $T_{h}<2T_{c}$ regime. As depicted by the red dashed lines, the efficiency consistently and strictly remains below the Carnot limit, demonstrating the absolute physical robustness of our cost-inclusive approach across all explored parameter spaces.
The fundamental reason we cannot provide a universal analytical proof for $\eta_{C}-\eta\geq 0$ under this specific $W_{\mathrm{cost}}$ bound lies in the structural depth of the theoretical framework itself. The analysis of the hybrid working substance-regenerator system is conducted exclusively through the reduced state of the working substance, and the operational cost of the regenerator is introduced ad hoc as a macroscopic penalty based on nature’s absolute ideal limit—the Carnot heat-pump bound. Specifically, this macroscopic limit assumes zero internal entropy production in the regenerator. On the other hand, there is an inevitable, positive entropy production inherently originating from the irreversible isochoric thermalization of the working substance; therefore, the Carnot efficiency naturally stands as the strict upper limit. Nature inherently protects its fundamental laws: the physical reality of the second law dictates that the actual regenerative mechanism must strictly demand a higher energetic cost ($W_{\mathrm{cost}}$) than the ideal, reversible Carnot limit precisely to compensate for this localized entropy production. This approach aligns perfectly with established methodologies in the literature; for instance, in hybrid quantum Otto cycle analyses involving system-quantum reservoir interactions, thermodynamic consistency is guaranteed by imposing an exact working substance parameter-dependent penalty term on the heat exchanges to enforce the Carnot bound [4, 40, 41]. Following this exact logic, instead of attempting to force a general proof from the purely macroscopic Carnot limit, we use Eq. (15) in Appendix B to provide a sufficient, parameter-dependent lower bound on $W_{\mathrm{cost}}$ that mathematically guarantees $\eta_{C}-\eta\geq 0$ across all regimes. Ultimately, the second law imposes a rigorous master condition, dictating that the true regeneration cost must satisfy the maximum of both the macroscopic thermodynamic constraint and the microscopic bound (see Appendix B for the full derivation details):
@@
As discussed, the current analytical bottleneck in universally proving the Carnot limit arises from treating the regenerator phenomenologically as a macroscopic reservoir. Moving beyond this hybrid approach and modeling the regenerator as an active quantum component will naturally resolve this mathematical conflict, allowing the Carnot bound to be analytically verified across all regimes. Therefore, to achieve a fully consistent and physically realistic framework in future studies, the system-regenerator coupling and correlations must be explicitly included. To this end, three candidate quantum-regenerator models suggest themselves: First, the regenerator could be modeled not as a Markovian heat bath but as a non-Markovian reservoir [41, 62, 63]. Unlike memoryless Markovian baths, non-Markovian reservoirs do not necessarily cause irreversible information loss and can exhibit information and energy back-flow [63], allowing a portion of the heat absorbed at stage (2) to be returned. Second, the regenerator can be modeled as an auxiliary quantum system on the same footing as the working substance [5, 47, 64, 65, 66, 54]. The heat rejected during the stroke (2) can induce eigenstate transitions in this auxiliary system, thereby storing energy in its populations. Algorithmic or adiabatic control protocols can subsequently be applied to reconvert the stored energy and regenerate the heat during the stage (4). Third, the regenerator may be described within a repeated–interaction (collision–model) framework, in which the working medium successively interacts with a finite chain of ancillary quantum systems that play the role of a structured reservoir with memory [67, 68, 38, 39]. During stage (2), a subset of these ancillas absorbs energy from the working medium, while during stage (4) the same ancillas are brought back into contact with the working medium and partially release the stored energy, thereby allowing partial heat regeneration. In all of the above proposals, the quantum modeling of the regenerator is intrinsically tied to the specific structure of the working substance; therefore, it is generally not possible to formulate a universal, model-independent quantum-regenerator description. Moreover, the isochoric strokes can terminate in non-equilibrium states. Consequently, additional thermalization steps are required prior to the subsequent isothermal contacts, so that the full cycle can be naturally described as a modified six-stroke regenerative Stirling cycle. Given that formulating a fully quantum regenerator model within any of these frameworks requires developing a dedicated, complex open-system architecture—each of which constitutes a standalone research problem in its own right—a rigorous mathematical and thermodynamic investigation of these possibilities is well beyond the scope of the present study and is therefore left for future work.